PREX/CREX Nov 14, 2019 14:00

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  373 507 581
 Bluejeans link:
 Offline room:   CEBAF Center F224-225


  1. Ongoing CREX prep activities: CREX-Install-Album
    1. Accelerator schedule
      • DJ: delayed by at least a week. Found a leak in bellows of girder 10. Needed to warm up 3 girders. Spin dance the weekend of Thanksgiving (highly speculative). CP: they are still discussing about potentially needing staffing for the halls on the week of Nov 24th.
      • CG: is hall C ready to start with their 1pass work? Are we going to be delayed? KP: not sure yet.
    2. S3 installation and electronics modification to incorporate the new S3 trigger.
      • ChG: installed two S3 scintillators on the left arm. Working on the logic triggers (summed up linear discriminated signal from the 3 paddles and made a logic signal from that). KP: this could give us some bias from the different sizes in the different paddles.
    3. Silviu: target
      • SC: target cooled down on Monday with no problems. Water is flowing and the vacuum is in mid 1e-7. Both ladders are home and locked.
      • KK: what do we know about the target alignment? SC: the preliminary results are on the docDB432. KP: we are aligned horizontally and 2 mm ofset vertically which is expected to be corrected after pulling vacuum. KK: has it been exercised to see what reproducibility? SC: everything seems to be ok.
    4. Kent: detector installation
      • KP: SAMs and detectors installed. Do we have integrating pedestals? Do we have counting pedestals? DA: will connect PMTs soon (the old ones; we will explore using new PMTs). KK: we should take some overnight runs to see how everything looks like and play the same games we played before PREX2 with updated (CREX) HVs. DM: we should run cosmics with the VDCs and the MD. We will deinstall the small US GEM by the end of the week. We will use the 5mm quartz we used during PREX2 and the same pre-amp setup.
    5. Alexandre: Compton e-det
      • AC: we need to have copies of all the signals we will need for the analysis. Juan Carlos and Adam will help with that.


In F245

Adam Z, Weibin Z, Juan Carlos C, Paul K, Robert R, Tao Y


Eric K, KK, Ciprian G, Dmitrii N, Ezekiel W, Alexandre C, Silviu CD, Caryn P, Ryan R, Chandan G, Devi A, Dustin M, Iris H, Don J, Kent P, Cameron C, Guido U, Victoria O

Cannot make it

Sanghwa P, Jinlong Z,